OpenStudy (unklerhaukus):

find \[\langle \theta \rangle \] for \(0≤\theta≤\pi/2\)

OpenStudy (unklerhaukus):

\[\langle x\rangle=\int\limits_0^{\pi/2}\theta\cdot\rho(\theta)\cdot\text d\theta \]

hartnn (hartnn):

whats \(\rho(\theta)\) ?

OpenStudy (unklerhaukus):

probability density

OpenStudy (unklerhaukus):

\[1=\int\limits_0^{\pi/2}\rho(\theta)\cdot\text d\theta\]

OpenStudy (unklerhaukus):

assuming a unifirm probability distribution \[\rho(\theta)=A\] \[\frac 1A=\int\limits_0^{\pi/2}\text d\theta\]

hartnn (hartnn):

A=2/\(\pi\) then

OpenStudy (unklerhaukus):

\[\frac 1A=\theta|_0^{\pi/2}=\frac\pi2\] \[A=\frac 2\pi\]

OpenStudy (unklerhaukus):

so, \[\langle x\rangle=\frac 2\pi\int\limits_0^{\pi/2}\theta\cdot\text d\theta\]

hartnn (hartnn):

pi/4 ?

OpenStudy (unklerhaukus):

\[=\frac 2\pi \frac{\theta^2}{2} |_0^{\pi/2}\] \[=\frac 2\pi \frac{\left(\pi/2\right)^2}{2}\] \[=\frac\pi 4\]

hartnn (hartnn):

nice, pdf was assumed or given? assuming it as uniform simplified it a lot.....

OpenStudy (unklerhaukus):

i probably should have stated uniform probability distribution in the question

OpenStudy (unklerhaukus):

so for \(0≤\theta≤2\pi\)\[\langle \theta \rangle =\pi\] the expected angle in a circle is 180°

OpenStudy (unklerhaukus):

makes sense

hartnn (hartnn):

uniform pdf will always give you mid-point as expectation....

hartnn (hartnn):

right ?

OpenStudy (unklerhaukus):

right